### v1

@@ Abstract
The classical belief revision framework proposed by Alchourron, Gardenfors, and Makinson involves the revision of a theory based on eight postulates. This paper focuses on exploring the revision theory based on quantum mechanics, known as the natural revision theory. In quantum systems, there are two reasoning modes: static intuitionistic reasoning, which incorporates contextuality, and dynamic reasoning, which is achieved through projection measurement.
@@
We combine the advantages of the two intuitionistic quantum logics proposed by Doering and Coecke respectively. We aim to provide a truth-value assignment for intuitionistic quantum logic that not only aligns with the characteristics of quantum mechanics but also allows for truth-value reasoning. We investigate the natural revision theory based on this approach.
@@
We introduce two types of revision operators corresponding to the two reasoning modes in quantum systems: object-level revision and operator-level revision, and we highlight the distinctions between these two operators. Unlike classical revision, we consider the revision of consequence relations in intuitionistic quantum logic. We demonstrate that, within the framework of the natural revision theory, both types of revision operators work together on the reasoning system of consequence relations. The outcomes of revision process are influenced by the order in which the interweaved operators are applied.
keywords:
@@
intuitionistic quantum logic , contextuality , projection measurement , truth-value reasoning , natural revision
journal: Annals of Pure and Applied Logic
[…]
[l1]organization=School of Mathematical Science, Beihang University, addressline=37 Xueyuan Road, Haidian District, city=Beijing, postcode=100191, country=China
1 Introduction
Classical belief revision was first proposed by Alchourron, Gardenfors, and Makinson in 1985[1]. A theory is contracted or revised in terms of six elementary and two supplementary postulates. They mainly researched how to appropriately select the subset consistent with the new information in the original theory according to the new one. Such selection criteria are usually subjective. In this paper, we mainly study the revision theory called natural revision, based on the objective laws of quantum mechanics.
@@ 1 Introduction
The logical characterization of quantum mechanics was first introduced by Birkhoff and von Neumann in 1936[2]. They showed that a quantum system can be described via Hilbert space. There is a one-to-one correspondence between projection operators of a quantum system and closed subspaces of Hilbert space. The operations between closed subspaces satisfy orthomodular law. Quantum logic represented by closed subspaces forming an orthomodular lattice is called standard quantum logic. It is still a hot spot in the field of quantum computing.
@@
The reasoning of quantum logic is non-monotonic reasoning based on features of quantum mechanics. Monotonicity means that new information added to the reasoning does not affect the result of the implication of old information. In a quantum system, when reasoning is revised by projection measurement as new information, some results implied from the old ones are no longer valid because the projection measurement changes the state of the system. We intend to investigate the nature of reasoning in quantum logic and thus provide a theory of natural revision determined by the objective laws of quantum mechanics.
@@
Standard quantum logic takes closed subspaces of Hilbert space as elements, the intersection of closed subspaces as the meet operation, the span of closed subspaces as the join operation, and the orthogonal complement of a closed subspace as the negation operation, which forms an orthomodular lattice called the property lattice. There is no implicative connective in the property lattice, so truth-value reasoning cannot be carried out.
@@
P.D. Finch[3, 4] defined the Sasaki hook $\rightsquigarrow_{s}$ and it is used as the implicative connective in property lattice. The left adjoint of Sasaki hook, called Sasaki projection $\wedge_{s}$, characterizes projection measurements in a quantum system. Roman and Zuazua[5] discussed that the Sasaki adjoint is some kind of deduction theorem. Engesser and Gabbay[6] inspected the link between non-monotonic consequence relations and belief revision. They argued that reasoning in quantum logic is a non-monotonic consequence relation, and the Sasaki projection is the revision operation of consequence relations. The adjoint operation Sasaki hook is the internalizing connective. A consequence revision system (CRS) is constructed by using formulas as operators. They examined the consequence relations based on truth-value reasoning in the property lattice, and regard $a,b,c$ in $a\wedge_{s}b\leq c$ as elements from the same level. With $\wedge_{s}$ as the revision operation, formulas $a$ and $b$ can be used as operators to revise a consequence relation. Sasaki adjoint can operate in the form of modus ponens, i.e., $a\wedge_{s}(a\rightsquigarrow_{s}b)\leq b$. However, because of the different meanings of elements in the operation, Sasaki adjoint cannot play the role of truth-value reasoning.
@@
Bob Coecke[7] indicated that Sasaki hook is not a static implicative connective but induces a dynamic one with a parameter in dynamic operational quantum logic. Define $\varphi_{a}^{*}(b):=a\wedge_{s}b$ as well as $\varphi_{a,*}(b):=a\rightsquigarrow_{s}b$, and $\varphi_{a}^{*}(b)\leq c\Leftrightarrow b\leq\varphi_{a,*}(c)\Leftrightarrow b\stackrel{{\scriptstyle\varphi_{a}}}{{\rightsquigarrow}}c$ is satisfied, in which $b$ and $c$ are variables while $a$ is a parameter. Obviously, they are elements from different levels. $\varphi_{a}^{*}(b)\leq c$ represents that with space $a$ as the projection space, the result of projecting $b$ space to $a$ is contained in $c$ space. In [6], considering only the revision form based on the projection operation is not enough to characterize quantum reasoning. We study the natural revision theory based on quantum logic by investigating the truth-value reasoning of quantum logic.
@@
In our study, we consider the application of topos theory for the truth assignment of properties in quantum systems. Isham et al.[8, 9] characterized physical quantities of a quantum system by presheaf structure. They emphasized that contextuality is a crucial characteristic of quantum systems, and properties in a quantum system should be assigned local truth values within specific contexts. Doering et al.[10, 11] defined a mapping called daseinisation to approximate a property across all contexts. They suggested that the global truth value of properties can be determined based on the outer daseinisation. Doering[12, 13] showed that elements corresponding to the outer daseinisations of properties generate a bi-Heyting algebra.
@@
However, upon investigating the implicative connective in this bi-Heyting algebra, we find that the implicative connective in the sense of Heyting algebra is too trivial, while the one in the sense of co-Heyting algebra is appropriately defined. Therefore, the elements selected by Doering constitute a co-Heyting algebra that satisfies the law of excluded middle, whereas ordinary truth-value reasoning should rely on a Heyting algebra that satisfies the law of non-contradiction as the reasoning structure.
@@
Coecke[14] constructed an intuitionistic quantum logic based on Heyting algebra using the downsets of elements in the property lattice. However, Coecke’s structure is only consistent with ordinary truth-value reasoning, which does not involve contextuality, and is insufficient to describe the characteristics of quantum mechanics. Thus, we propose an intuitionistic quantum logic that combines the respective advantages of the two intuitionistic quantum logics mentioned above. This logic satisfies the characteristics of quantum mechanics and allows for truth-value reasoning.
@@
In our approach, we propose selecting elements according to the inner daseinisation for the global truth value of properties. This selection ensures that the chosen elements constitute a Heyting algebra. Truth-value reasoning within a Heyting algebra is static reasoning. However, projection measurement, as an external action, transforms one Heyting algebra into another, and the reasoning generated by the Sasaki adjoint represents dynamic truth-value reasoning between Heyting algebras.
@@
We believe that a consequence relation occurs within a Heyting algebra. There are two types of revision operations for a consequence relation. The first is to add new assumptions into the Heyting algebra as new information, corresponding to the $\wedge$ operation as revision, which we refer to as object-level revision. The second is the action of projection measurement between Heyting algebras as new information, corresponding to the $\wedge_{s}$ operation as revision, which we refer to as operator-level revision. Both types of revision use formulas as operators to revise a consequence relation. The difference lies in the level of the revision operator and the method of revising the antecedents of a consequence relation. Object-level revision changes the consequence relation within a Heyting algebra to another one within the same Heyting algebra, while operator-level revision changes the consequence relation within a Heyting algebra to one within another Heyting algebra.
@@
Engesser et al. utilized the orthomodular lattice to represent the natural revision theory. However, the orthomodular lattice fails to differentiate between the two levels of operations, resulting in a single reasoning mode. Consequently, their work [6] only includes one revision operator, which corresponds to the “expansion” in the revision theory. In contrast, intuitionistic quantum logic allows for the distinction between static reasoning and dynamic one. In our approach, we propose two levels of revision operators that correspond to the “expansion” and “revision” in the revision theory, respectively. The outcomes of revision process are influenced by the order in which the interweaved operators are applied. These operators enable the description of the natural revision theory following the fundamental characteristics of quantum mechanics.
@@
In section 2, we provide the basic knowledge and background relevant to this paper. In section 3, we review the existing revision theory, discuss the revision of consequence relation based on the operations in the orthomodular lattice by Engesser et al., and highlight the shortcomings of this method. We then investigate the truth assignment of properties in a quantum system based on topos theory, analyze the role of the law of excluded middle and the law of non-contradiction in the structure proposed by Doering et al., and present a more suitable method for truth-value reasoning in section 4. In section 5, we discuss the static and dynamic reasoning of quantum logic based on the provided truth assignment, consider the two types of revision operators for a consequence relation, and present the natural revision theory based on the truth-value reasoning of quantum logic. Finally, we summarize the paper and propose prospects for research.
2 Preliminaries
2.1 Standard quantum logic
@@
Given a Hilbert space $\mathcal{H}$, denote a collection of its closed subspaces $Sub(\mathcal{H})$. Define partial relation $\leq$ to be the inclusion of closed subspaces. $\wedge$ operation is the intersection of two closed subspaces. $\vee$ is the span of two closed subspaces. $\neg$ is the complement of a closed subspace. $\langle Sub(\mathcal{H}),\leq,\wedge,\vee,\neg\rangle$ is an orthomodular lattice. There is a one-to-one correspondence between closed subspaces of a Hilbert space and projectors in a quantum system, and a projector represents a property in the system. Hence the orthomodular lattice is called a property lattice. Without causing ambiguity, we denote the property lattice by $L$.
Definition 3.
@@
The internal operations of the property lattice are not sufficient to characterize quantum mechanics, so a pair of external operations $\wedge_{s}$ and $\rightsquigarrow_{s}$ are introduced into the property lattice. Furthermore, the mappings $a\wedge_{s}\_$ and $a\rightsquigarrow_{s}\_$ form a pair of adjoints as well, called Sasaki adjoint, which is denoted by $a\wedge_{s}\_\dashv a\rightsquigarrow_{s}\_$. In which
@@
is analogous to Modus Ponens in logical reasoning mechanism and to obtain
@@
Introducing Sasaki adjoint into property lattice $L$ as a mathematical tool to characterize quantum logic, which is called standard quantum logic.
2.2 Topos quantum theory
@@
Definition 4.
A context is a n-dimensional commutative von Neumann subalgebra $V$. The set of generators of $V$ is denoted by $\mathcal{F}_{V}=\{\hat{P_{1}},\hat{P_{2}},...,\hat{P_{n}}\}$. Every component of the spectral decomposition of self-adjoint operators in $V$ is an element in $\mathcal{F}_{V}$. Every self-adjoint operator in $V$ is a linear combination of the generators above. If we only consider all projectors in $V$, which is denoted by $P(V)$, then $P(V)$ constitutes a Boolean subalgebra.
Definition 5.
The set $\sigma(V)=\{\lambda_{1},\lambda_{2},...,\lambda_{n}\}$ is defined to be the spectrum of an n-dimensional commutative von Neumann algebra $V$, where every $\lambda_{i}$ maps a self-adjoint operator in $V$ to the coefficient of $\hat{P_{i}}$ component of spectral decomposition of the operator. $\lambda_{i}$ is of one-to-one correspondence to $\hat{P_{i}}$.
Definition 6.
@@
A property does not belong to every context in general. Doering[10, 11] indicates that for the projector corresponding to a property, one can find an approximation of the projector in every context.
Definition 8.
@@
Notice that Doering uses only the outer daseinisation when approximating a projector via daseinisation. Without causing confusion, Doering denotes outer daseinisation by daseinisation briefly. However, in this paper, we need to distinguish outer and inner daseinisation.
2.3 Consequence relation
@@
$A\vdash_{\Delta}B\qquad iff\qquad B\in\Delta\circ A.$
@@
3 Revision theory in standard quantum logic
In their work on standard quantum logic, Engesser et al.[6] propose a revision theory based on the Sasaki adjoint operation of property lattice and treat all the elements in the operation indistinctively. However, Bob Coecke[7] showed that the roles of elements in Sasaki adjoint are different. Hence the revision theory defined by standard quantum logic may not conform to the nature of quantum mechanics and not reflect its features precisely.
Definition 9.

### v2

@@ Abstract.
The classical belief revision framework, as proposed by Alchourron, Gardenfors, and Makinson, involves the revision of a theory based on eight postulates. In this paper, we focus on the exploration of a revision theory grounded in quantum mechanics, referred to as the natural revision theory.
@@
There are two reasoning modes in quantum systems: static intuitionistic reasoning, which incorporates contextuality, and dynamic reasoning, which is achieved through projection measurement. We combine the advantages of two intuitionistic quantum logic frameworks, as proposed by Döring and Coecke, respectively. Our goal is to establish a truth-value assignment for intuitionistic quantum logic that not only aligns with the inherent characteristics of quantum mechanics but also supports truth-value reasoning. The natural revision theory is then investigated based on this approach.
@@
We introduce two types of revision operators that correspond to the two reasoning modes in quantum systems: static and dynamic revision. Furthermore, we highlight the distinctions between these two operators. Shifting away from classical revision paradigms, we consider the revision of consequence relations in intuitionistic quantum logic. We demonstrate how, within the natural revision theory framework, both revision operators collectively influence the consequence relations. Notably, the outcomes of revision process are impacted by the sequence in which these interweaved operators are deployed.
Key words and phrases:
@@
intuitionistic quantum logic, contextuality, projection measurement, truth-value reasoning, revision operator
1991 Mathematics Subject Classification:
Primary 81P10, 81P13, 81Q10, 81Q35, 03G12; Secondary 06C15, 06D20
1. Introduction
Classical belief revision was first proposed by Alchourron, Gardenfors, and Makinson in 1985[1]. A theory is contracted or revised in terms of six elementary and two supplementary postulates. Their research centered on the methodology for selecting subsets from an original theory that align with newly acquired message, a process typically influenced by subjective criteria. In this paper, our primary focus is on the exploration of a revision theory known as “natural revision,” which is grounded in the fundamental principles of quantum mechanics. This methodology enables a selection process governed by objective criteria, thereby distinguishing it from methods reliant on subjective judgment.
@@ 1. Introduction
The logical characterization of quantum mechanics was introduced by Birkhoff and von Neumann in 1936[3]. They showed that a quantum system can be described via Hilbert space. There is a one-to-one correspondence between projection operators of a quantum system and closed subspaces of Hilbert space. The operations between closed subspaces satisfy orthomodular law. Quantum logic, represented by an orthomodular lattice composed of closed subspaces, is referred to as standard quantum logic. It is still a hot spot in the field of quantum computing[24, 2].
@@
The reasoning of quantum logic is non-monotonic reasoning based on features of quantum mechanics. Monotonicity means that new message added to the reasoning does not affect the result of the implication of old messages. In a quantum system, when reasoning is revised by projection measurement as new message, some results implied from the old ones are no longer valid because the projection measurement changes the state of the system. We intend to investigate the nature of reasoning in quantum logic and thus provide a natural revision theory determined by the fundamental laws of quantum mechanics.
@@
Standard quantum logic takes closed subspaces of Hilbert space as elements, the intersection of closed subspaces as the meet operation, the span of closed subspaces as the join operation, and the orthogonal complement of a closed subspace as the negation operation, which forms an orthomodular lattice called the property lattice. There is no implication connective in the property lattice, so truth-value reasoning cannot be carried out.
@@
Finch[13, 14] defined the Sasaki hook $\rightsquigarrow_{s}$ and it is used as the implication connective in property lattice. The left adjoint of Sasaki hook, called Sasaki projection $\wedge_{s}$, characterizes projection measurements in a quantum system. Roman and Zuazua[23] discussed that the Sasaki adjoint is some kind of deduction theorem. Engesser and Gabbay[12] inspected the link between non-monotonic consequence relations and belief revision. They argued that reasoning in quantum logic is a non-monotonic consequence relation, and the Sasaki projection is the revision operation of consequence relations. The adjoint operation Sasaki hook is the internalizing connective. A consequence revision system (CRS) is constructed by using formulas as operators. They examined the consequence relations based on truth-value reasoning in the property lattice, and regard $a,b,c$ in $a\wedge_{s}b\leq c$ as arguments. With $\wedge_{s}$ as the revision operation, formulas $a$ and $b$ can be used as operators to revise a consequence relation. Sasaki adjoint can operate in the form of modus ponens, i.e., $a\wedge_{s}(a\rightsquigarrow_{s}b)\leq b$. However, because of the different meanings of elements in the operation, Sasaki adjoint cannot play the role of truth-value reasoning.
@@
Bob Coecke[6] indicated that Sasaki hook is not a static implication connective but induces a dynamic one with a parameter in dynamic operational quantum logic. Define $\varphi_{a}^{*}(b):=a\wedge_{s}b$ as well as $\varphi_{a,*}(b):=a\rightsquigarrow_{s}b$, and $\varphi_{a}^{*}(b)\leq c\Leftrightarrow b\leq\varphi_{a,*}(c)\Leftrightarrow b\stackrel{{\scriptstyle\varphi_{a}}}{{\rightsquigarrow}}c$ is satisfied, in which $b$ and $c$ are variables while $a$ is a parameter. Obviously, the meaning of $a$ in the operation differs from that of $b$ and $c$. $\varphi_{a}^{*}(b)\leq c$ represents that with space $a$ as the projection space, the result of projecting $b$ space to $a$ is contained in $c$ space. In [12], considering only the revision form based on the projection operation is not enough to characterize quantum reasoning. We investigate the natural revision theory based on quantum logic by exploring both static and dynamic aspects of truth-value reasoning in quantum logic.
@@
In our study, we explore the utilization of topos theory to depict the contextuality of properties within quantum systems. Isham et al.[18, 17] characterized physical quantities of a quantum system by presheaf structure. They emphasized that contextuality is a crucial characteristic of quantum systems, and properties in a quantum system should be assigned local truth values within specific contexts. Döring et al.[10, 11] defined a mapping called daseinisation to approximate a property across all contexts. They suggested that the global truth value of properties can be determined based on the outer daseinisation. Döring[8, 9] showed that elements corresponding to the outer daseinisations of properties generate a bi-Heyting algebra.
@@
However, upon investigating the implication connective in this bi-Heyting algebra, we find that the implication connective in the sense of Heyting algebra is trivial, while the one in the sense of co-Heyting algebra is non-trivial. Therefore, the elements selected by Döring constitute a co-Heyting algebra that satisfies the law of excluded middle, whereas ordinary truth-value reasoning should rely on a Heyting algebra that satisfies the law of non-contradiction as the reasoning structure.
@@
Coecke[4] constructed an intuitionistic quantum logic based on Heyting algebra using the downsets of elements in the property lattice. However, while Coecke’s structure aligns with conventional truth-value reasoning, it does not involve contextuality, and is insufficient to describe the characteristics of quantum mechanics. In their research on the dynamic aspect of quantum logic[5], Coecke et al. highlighted that a causal relationship based on actions can be induced to replace the partial order relationship between elements. They defined a pair of adjoint according to this causal relationship, where the adjoint operations respectively represent the propagation and causation of properties. Reasoning in dynamic operational quantum logic is then conducted based on this adjoint relationship. We intend to propose an intuitionistic quantum logic that combines the respective advantages of static and dynamic quantum logics mentioned above. This logic satisfies the characteristics of quantum mechanics and allows for truth-value reasoning.
@@
In our approach, we propose selecting elements according to the inner daseinisation for the global truth value of properties. This selection guarantees that the chosen elements form a Heyting algebra. Truth-value reasoning within a Heyting algebra is static reasoning. However, projection measurement, serving as an external action, transforms one Heyting algebra into another, and the truth-value assignment generated by the Sasaki projection represents dynamic truth-value changes between Heyting algebras.
@@
We believe that a consequence relation occurs within a Heyting algebra. There are two types of revision operations for a consequence relation. The first is to add new assumptions into the Heyting algebra as new message, corresponding to the $\wedge$ operation as revision, which we refer to as static revision. The second is the action of projection measurement between Heyting algebras as new message, corresponding to the $\wedge_{s}$ operation as revision, which we refer to as dynamic revision. Both types of revision use formulas as operators to revise a consequence relation. The distinction lies in the methodology employed to revise the antecedents of a consequence relation. Static revision changes the consequence relation within a single Heyting algebra, resulting in a new consequence relation that remains within the original Heyting algebra. Meanwhile, dynamic revision involves altering the consequence relation from one within a given Heyting algebra to a new consequence relation situated within a different Heyting algebra.
@@
Engesser et al. utilized the orthomodular lattice to represent the natural revision theory. However, the orthomodular lattice fails to differentiate between the static and dynamic operations, resulting in a single reasoning mode. Consequently, their work [12] only includes one revision operator, which corresponds to the “expansion” in the revision theory. In contrast, intuitionistic quantum logic allows for the distinction between static reasoning and dynamic one. In our approach, we propose two kinds of revision operators that correspond to the “expansion” and “revision” in the revision theory, respectively. The outcomes of revision process are influenced by the order in which the interweaved operators are applied. These operators enable the description of the natural revision theory following the fundamental characteristics of quantum mechanics.
@@
In section 2, we provide the basic knowledge and background relevant to this paper. In section 3, we review the existing revision theory, discuss the revision of consequence relation based on the operations in the orthomodular lattice by Engesser et al., and highlight the shortcomings of this method. We then investigate the truth assignment of properties in a quantum system based on topos theory, and present a more suitable method for truth-value reasoning in section 4. In section 5, we discuss the static and dynamic reasoning of quantum logic based on the provided truth assignment, consider the two types of revision operators for a consequence relation, and present the natural revision theory based on the truth-value reasoning of quantum logic. Finally, we summarize the paper and propose prospects for research.
2. Preliminaries
2.1. Standard quantum logic
@@
Given a Hilbert space $\mathcal{H}$, denote a collection of its closed subspaces $Sub(\mathcal{H})$. Define partial order relation $\leq$ to be the inclusion of closed subspaces. $\wedge$ operation is the intersection of two closed subspaces. $\vee$ is the span of two closed subspaces. $\neg$ is the complement of a closed subspace. $\langle Sub(\mathcal{H}),\leq,\wedge,\vee,\neg\rangle$ is an orthomodular lattice. There is a one-to-one correspondence between closed subspaces of a Hilbert space and projectors in a quantum system, and a projector represents a property in the system. Hence the orthomodular lattice is called a property lattice. Without causing ambiguity, we denote the property lattice by $L$.
Definition 2.3.
@@
A pair of operations $\wedge_{s}$ and $\rightsquigarrow_{s}$ are introduced into the property lattice to describe the characteristics of quantum mechanics. Furthermore, the mappings $a\wedge_{s}\_$ and $a\rightsquigarrow_{s}\_$ form a pair of adjoints as well, called Sasaki adjoint, which is denoted by $a\wedge_{s}\_\dashv a\rightsquigarrow_{s}\_$. In which
@@
is analogous to modus ponens in formal systems and to obtain
@@
2.2. Topos quantum theory
@@
Definition 2.4.
A context is a commutative von Neumann subalgebra. For a $n$-dimensional context $V$, the set of generators of $V$ is denoted by $\mathcal{F}_{V}=\{\hat{P_{1}},\hat{P_{2}},...,\hat{P_{n}}\}$. Every component of the spectral decomposition of self-adjoint operators in $V$ is an element in $\mathcal{F}_{V}$. Every self-adjoint operator in $V$ is a linear combination of the generators above. If we only consider all projectors in $V$, which is denoted by $P(V)$, then $P(V)$ constitutes a Boolean subalgebra.
Definition 2.5.
Given a $n$-dimensional commutative von Neumann algebra $V$, define the set $\sigma(V)=\{\lambda_{1},\lambda_{2},...,\lambda_{n}\}$ to be the spectrum of $V$, where every $\lambda_{i}$ maps a self-adjoint operator in $V$ to the coefficient of $\hat{P_{i}}$ component of spectral decomposition of the operator. $\lambda_{i}$ is of one-to-one correspondence to $\hat{P_{i}}$.
Definition 2.6.
@@
A property does not belong to every context in general. Döring[10, 11] indicates that for the projector corresponding to a property, one can find an approximation of the projector in every context.
Definition 2.8.
@@
It should be noted that Döring employs only the outer daseinisation for approximating a projector. Without causing confusion, Döring refers to outer daseinisation simply as daseinisation. However, in this paper, it is necessary to make a clear distinction between outer and inner daseinisation.
2.3. Consequence relation
@@
$A\vdash_{\Delta}B\qquad\text{iff}\qquad B\in\Delta\circ A.$
@@
3. Revision theory in standard quantum logic
In their work on standard quantum logic, Engesser et al.[12] propose a revision theory based on the Sasaki adjoint operation of property lattice, treating all elements within this operation uniformly. However, Bob Coecke[6] has demonstrated that the roles of elements in Sasaki adjoint are distinct. This discrepancy suggests that the revision theory defined by standard quantum logic may not fully align with the fundamental law of quantum mechanics, nor accurately capture its characteristics.
Definition 3.1.